Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

Key: BOLD: Minutes/notes from last week; BLUE: Notes from this week.

News

  • OSRO: identified issues mostly resolved; RSRO Science observing begun! More details below; special expanded RSRO section.
  • Staff Travel:
    • Lazio: 28 Jun - 06 Jul
    • Butler: 01 Jul - 11 Jul
    • Clark: 04 Jul - 11 Jul
    • Krauss: 23 Jul - 27 Jul
    • McMullin: 26 Jul - 30 Jul
    • van Moorsel: 26 Jul - 30 Jul
  • Visitors:
    • Brent Carlson: 5 July- 16 July
  • Key Dates:
    • PASEO Meeting: 15-16 July; Walk-throughs this week.

Critical Commissioning review/reports

  • Overall status: Several updates to CBE software; believed to have resolved CBE node failure problem (Robnett/Rupen).
  • Highest priority development areas (Note - changing this week.):
    • 1. OSRO blockers: Bottom line is vast improvement here (we've gone from a chaotic zoo of phenomena to a fairly well understood set of situations that have been greatly mitigated). Considered unblocked at this stage though residual efforts are still underway.
      • Details:
        • Phase jumps: additional cables have been added to the two worst antennas (my understanding is that when we get close to the edge of a range, we can get the jumps, the addition/subtraction of cables moves it into the middle of the range and things track smoothly); this has resolved the problem. We didn't do this when we transitioned to 20 Hz (from 19.2) thinking that updates in the software removed the necessity. This was not the case. For now, we'll fix up the other egregious antennas to remove the phase jumps and revisit the software at a later date.
        • Amplitude jumps: the rapid monitoring exposed that set-and-remember wasn't working properly on several antennas; I don't think we've done the complete census but EA17 came into the array after these checks (and had a different down-converter than the rest); others there are suspicions of replacements but the situation is still clarifying. Regardless, these are fixable once identified and this is going on (1dB jumps only).
        • Loss of coherence: 2 styles: 1) 14s before the end of scan prior to a band change; found that the side band flippers were prematurely firing; looks like the executor and this should be readily solved. 2) 10s at the beginning of a scan seems to be a late application of the SSLOs (also likely the executor but stay tuned).
      • Remaining issues:
        • Missing scans
    • 2. Wide band widths (2 GHz)
      • Tests (1 hour trials) taken on several more projects. First science observing (See RSRO section). Restrictions due to failures in the CBE nodes/NICs; CBE node failure was explored down several fronts (different Lustre software, overdrawing on power, Red Hat kernal bugs); each addressed, with the last likely the ultimate culprit; running without failures for several days.
    • 3. Switched power implementation
      • Document delivered to SSS for implementation. to join the various data flow developers (StB->MCAF->SDM) in how to effectively implement this addition (in particular the mapping and flow of information as different parts of the system are not cognizant of things like IFs and scans). All issues seemed to be resolved; this work will go on so that by the time Rich returns from vacation, he can do the final development to put the switched power into the SDM.
    • 4. 3-bit sampler testing
      • Bottom line; working mostly in the background this has gone smoothly; 4 installed, 3 working (last needs an LO modification); fringes on all baselines (3 antennas) for the single IF pair. One complication is in the SDM where some agreed upon expanded enumerations (allowing the EVLA-only to distinguish between 3 and 8 bit sampler products) has not been implemented. It's being pushed on. Other signal path differences are being worked on which likely need to toggle depending on 3/8 bit situations.

  • Revised priorities.
    • 1) Commissioning tests/system/correlator stability (particularly focused on using Brent's time).
    • 2) RSRO science observing
    • 3) 3-bit sampler testing
    • 4) OSRO science observing
    • 5) Demo science observing (getting at least one project through is high priority; likely able to complete a subset of the 'high' priority; currently looking at: Cyg A, Crab Nebula, M87, M51, IRC+10216 first).

System Verification/Hardware Commissioning:

Testing and characterization of the array based on the design goals.

Targets

Area Target Priority JIRASorted ascending Deliver Status
* Switched power gain correction 0 - pending SDM implementation for further work  
* RFI Sweep for all bands 2 - 2010Q3 Pending wide band widths
      • Still need to characterize phase transfer; issues with Cyg A imaging.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Med Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; understood as a synchronization issue; currently at 5% of the data for 'most' observations. Carlson/Fort at work on mitigating/eliminating.

Reports

  • Antenna: Perley
    • EA06 status IF B module replaced; problem fixed.
    • EA20 amplitude modulations. fixed.
    • syscoll observation report. (Sowinski)
  • Receiver: Perley
    • EA12 Q band Rx installed and working (Sowinski).
    • Ku rxs coming (2) next week.
  • Array: Perley
    • Wideband testing reports.
      • X-band wide (Perley)
      • Reference pointing (Chandler)
      • RSRO Testing:
  • Correlator: Rupen
    • Status:
    • Succinct list of issues
      • *5% zeros/dropped data due to loss of synchronization; working to improve this (Carlson); 1% switching bands; few% switching configurations.
      • Missing scans
      • X band frequency labeling/sub-band ordering.

Observing Band Checkout

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

Targets

Reports

  • Low Frequency (L/S/C): Momjian
    • News/Data Updates: 10B-187 data review: Baselines to EA02 (E02) show a comb in Stokes LL throughout the observations (for instance, see the attached EA02_SPEC.png). The sky frequency is 1420.40575 MHz, and the bandwidth is 1 MHz. The spectral spikes are separated by 0.125 MHz. (ECDATA-261). Being reviewed by Kerry.
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-38

  • Mid Frequency (X/Ku): Dhawan
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links

  • High Frequency (K/Ka/Q): Sjouwerman/Claussen
    • Status:
      • Operating; users should include one of the canonical flux calibrators in SB
      • Reference pointing in C or X
      • No observing at Ka < 32 GHz; No longer blocked; SSS/OPT code is consistent with limitations in hardware.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-30

Observing Mode Support:

Testing and development within broad observing modes distinguished by their science goals (continuum, spectral line, polarization).

Targets

Area Target Priority JIRA Deliver Status
Spectral Line Bandpass stability 1 ECSV-35 2010Q3  
Spectral Line Doppler tracking 1 ECSV-32 2010Q3  
Spectral Line Spectral line setups; calibration procedures recommendations (OSRO/RSRO) 1 ECSV-31 2010Q3  
Polarization On-axis leakage calibration 1 ECSV-25 2010Q3  
Polarization R-L phase calibration 1 ECSV-25 2010Q3  
Polarization Observer Recommendations (OSRO/RSRO) 2 ECSV-25 2010Q3  
Polarization Beam squint calibration 2 ECSV-25 2010Q3  
Polarization Primary beam (off-axis) calibration 2 ECSV-25 2010Q3  
Polarization Polarization calibration service observation plan 3 ECSV-25 2010Q4  
Demo Science SBs for approved EPO projects 1 ECSV-57 2010Q3 Started for a subset

Reports

  • Spectral Line Modes (Ott/Goss): TSPE
    • Testing:
      • Spectral Line general commissioning tests (diff spw widths, diff frequencies, bandpasses, etc): ECSV-92 * Bandpass stability: ECSV-35: Report
    • Links: ECSV-19
  • Polarization Modes (Myers): TPOL
    • Status: Pending reports
    • Testing
      • Low frequency check outs completed; derived fractional polarization for known sources agree with predictions.
      • Multi-band leakage/pos. angle calibration report pending.
      • Post-processing (AIPS): AH1004 data checkout (still needed)
    • Links:

  • Demo Science (Chandler/Condon): TDEM
    • Status: Begun; list requires prioritization; limited remaining time in D-config
    • Testing
      • Cyg A (Perley): SBs ready; Issues as noted above; will need repeat.
      • Crab Nebula (Frail): SBs submitted; 1 hour trial report (Frail)
    • Links: https://staff.nrao.edu/wiki/bin/view/EVLA/DemoScience

  • Advanced Planetary: TPLA
    • Testing
      • Cassini observations obtained.; Status.
    • Status:
      • Executor: Ready for testing (Butler).
      • SDM: Ephemerides/motion not supported.
  • Advanced: Solar: TSOL
    • Status: Planning/prototype testing ongoing; working with L, Ku (special Rx setups in these two bands). Perley: On sun fringe check with WIDAR.
  • Advanced: Pulsar (Brisken): TPUL
  • Advanced: VLBI (Walker/Brisken): TVLB
  • Advanced: Astrometry (Mioduszewski)

Observing Support:

Testing and development of the end-to-end capability of the array to support observations to the community (Documentation, OPT, Scheduling, Instrument models, Calibrator survey/models DB, Rapid response observing, demo science, post-processing).

Targets

Area Target Priority JIRA Deliver Status
Software CASA development cycle priorities 0 -   Spreadsheet developed; time allocations needed for P1s
Software CBE averaging/decimation (time/freq) 1 - 2010Q3; vci averaging - honors other-than-1s dumps.  
Software/Calibration Recommended data path for weights in calibration 1 - 2010Q3  
Flux Calibration SBs for Flux calibrator models 1 ECSV-43 2010Q3 Needed for D config (2 freqs/band); 5 h of observations ready

Reports

  • Meta summary
    • Multi-config SBs in regular use; note 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

  • Software Systems (Overlord: Butler):
    • Versioning Page: https://staff.nrao.edu/wiki/bin/view/EVLA/EVLASoftwareVersioning Updated.
    • CBE store - new lustre file system next Wednesday. Happened yet
    • OPT (Lorant):
    • Scheduling (Joan):
      • In use (partially; many system tests that require more time management).
    • Executor (Barry):
    • Corr (CBE/CM; Michael): Sonja has departed.
    • Archive (John):
    • Post-processing
      • CASA (Nick):
        • Development plan out the door; being finalized then will post; most priority 1 development items covered; some speed improvements in flagging (being tested by Gustaaf/Joe); proposed change to flagdata interface; needs more thought (asked Miriam to lead a draft set of use cases to refine the interface options); problem if 'source' init casapy script (when adding own tasks) and running casapy - Wes is looking into it; discussion of timing of updates to CASA dependent libraries - can request updates to be judged by Nick. Finalized plan draft sent to developers - not clear time allocations. Needs plan of work.
      • AIPS (Eric): REWAY statistics
    • ASDM
      • Priority 1: SDM total power/switched power table write-up (Rupen). Done.
      • Priority 2: Flag table write-up (Myers/McMullin) Due. Needed for 5 Jul ALMA FBT
      • Priority 3: Ephemeris table resolution (Butler)

  • Flux Calibrator Models (Amy/Rick): TCAL
    • Report: TCAL0001_sb1576498_1.55377.02314818287, TCAL0001.sb1575783.eb1620753.55378.450005439816
    • Links: ECSV-43

  • RSRO/ECSO: TRSR
    • Status: RSRO Readiness (Cycle1) Master List
      • Crystal and Todd headed back to CV
    • RSRO Science Observations (Red=recent observations requiring report; Blue=change; Black=pending):
      • AB1345 (TRSR0004): TRSR0001_sb1620012_1_002.55380.22739263889 (Rupen)
      • AC992: AC992_sb1426692_1_000.55380.88644908565 (Chomiuk)
      • AB1346 (TRSR0005): AB1346.55382.19686146991 (Brogan/Hunter)
      • AS1015 (TRSR0010): AS1015_sb1626830_1.55380.92802219908 (Chomiuk)
      • AV318: Converted to OSRO and completed.
      • AM1014 (TRSR0009)
      • AT374 (TRSR0011)
      • AK726 (TRSR0007)
      • AA330 (TRSR0003)
      • AL746 (TRSR0008): TRSR0008_Ka.55382.59489054398, TRSR0008_C.55382.62113342593
      • AC982 (TRSR0002)
      • AH1006 (TRSR0006)
    • Resident Schedule: Posted on home page of twiki now!
      • 01 Jan - 16 May: Adam Deller (AC982: RSRO-4): 300A
      • 31 Jan - 31 Jul: Laura Chomiuk (AS1015/AK726: RSRO-11/RSRO-6): 200C
      • 04 Feb - 01 Jun: Jacqueline van Gorkom (AV318: RSRO-13): 200D
      • 25 Feb - 01 Jun: James Miller-Jones (AM1014: RSRO-9): 200E
      • 02 Mar - 07 Jun: Manuel Aravena (AA330: RSRO-1): 300G
      • 16 Mar - 08 Apr: Dave Green (AB1345: RSRO-2): 300G
      • 20 Apr - 30 Jun: Crystal Brogan (AB1346: RSRO-3): 200B
      • 20 Apr - 30 Jun: Todd Hunter (AB1346: RSRO-3): 200B
      • 30 Apr - 15 Jul: Joseph Lazio (AC982: RSRO-4): 266
      • 07 Jun - 16 Sep: Laura Perez (AC982: RSRO-4): 200E
      • 28 Jun - 05 Oct: Volker Heesen (AH1006: RSRO-5): 200G
      • 05 Jul - 26 Oct: Russ Taylor (AT374: RSRO-12): 300G
      • 21 Jul - 23 Sep: Dave Green (AB1345: RSRO-2): 300G
      • 15 Sep - 15 Dec: Adam Leroy (AL746: RSRO-7): 200C
      • 01 Oct - 01 Jan: Hendrik Linz (AC982: RSRO-4): 200C
      • 01 Jan - 01 Apr: Jane Greaves (AC982: RSRO-4): 300G
      • 02 Jan - 02 Apr: Stuart Corder (AC982: RSRO-4): 300G
    • Current Residents
      • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
        • This week:
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
        • This week:
      • Joe Lazio (POC: Claire Chandler): Low frequency observing.
        • This week:
      • Volker Heesen (POC: ):
        • This week:
    • Past (and future residents)
      • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
      • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
      • Manuel Aravena (POC: Chris Carilli); area(s): high frequency observing
      • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
      • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)

  • Science Support Team
    • McMullin: Scheduling/On-call; test data reviews; review X/C pointing results; PASEO Prep; FLAG table (with Steve)
    • Myers: Review of TPOL0002 with Crystal looking at cal transfer
    • Ott: CASA development planning; work for CDR 8; submit M51 SBs (demo science); helpdesk tickets.
    • Greisen: catch-up on e-mail; AIPS letter; change scaling on cubes (Jy/beam!).
    • Dhawan: data reviews from targeted tests, pointing (plots, reports).
    • Kogan: FRMAP task tutorial
    • Rupen: Switched power writeup; testing to isolate outstanding OSRO blockers; RSRO testing; work with Sonja.
    • Krauss: Reviewing pipelines; developing EVLA doc pages, improve organization, particularly for non-radio astronomers.

Advanced Science Support:

These are the testing/development areas needed to fulfill the full capabilities of the EVLA along with additional observing modes for other branches of the community.

  • Algorithms (Owen/Bhatnagar)

  • Astrometry (Mioduszewski)
    • Status: Wait for A array(?)

AOB

Misc Actions

  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 05 Jul 2010
Topic revision: r6 - 2010-07-06, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback